OpenStudy (anonymous):

need help.. 6th grade math fan and medal!

OpenStudy (anonymous):

Which of the values for x and y make the equation 2x + 3y + 4 = 15 true? x = 6, y = 3 x = 3, y = 5 x = 6, y = 4 x = 1, y = 3

RhondaSommer (rhondasommer):

what do you think?

OpenStudy (anonymous):

D ??

OpenStudy (lynfran):

start by substituting the given values into the equation to find out... when x=6 2x+3y+4=15 2(6)+3y+4=15 .. solve for y to find out ... do so for each given values of x"..

OpenStudy (anonymous):

how bout y??

RhondaSommer (rhondasommer):

so first subtract four from both sides of the equation

OpenStudy (anonymous):

well i did the math on all of them but i got higher numbers

RhondaSommer (rhondasommer):

okay...hold on

OpenStudy (lynfran):

D is correct when x=1 2(1)+3y+4=15 2+4+3y=15 6+3y=15 3y=15-6 3y=9 y=3

OpenStudy (anonymous):

oh ok so i was right then. thx for help!!

OpenStudy (anonymous):

i faced both of u girls

OpenStudy (anonymous):

faned

RhondaSommer (rhondasommer):

she didn't do it right :/ subtract four from both sides \[2x+3y=11\] subtract 2x from both sides \[3y=-2x+11\] divide 3 to get y alone \[y=(-2x+11)/3\] insert x (1) and plug in the answer is correct but she didn't simplify the equation like she was supposed to.